5grams of CaCO3 reacts completely with excess HCL for 2 minutes, find the rate of reaction in grand per hour
1
Expert's answer
2020-03-27T09:01:58-0400

"2 minutes (\frac{1 hour}{60 minutes}) =\frac{1}{30} h"


"Rate = \frac{\Delta m(CaCO_3)}{\Delta Time} = \frac{5 grams}{\frac{1}{30} h} =150 g/h"